# Fast-Convergent Proximity Graphs for Approximate Nearest Neighbor Search

### v4

@@ 7. Conclusions
    This paper introduces $\alpha$-CG, a new PG structure for high-performance ANN search. Specifically, $\alpha$-CG employs a well-designed pruning rule — shifted-scaled triangle inequality — to eliminate ineffective candidates. We prove that, under a realistic assumption that the distance between the query and its NN is bounded by a constant $\tau$, $\alpha$-CG guarantees exact NN retrieval in poly-logarithmic time. Without this assumption, it ensures ANN search within the same complexity bounds. To reduce graph construction overhead, we develop an approximate variant $\alpha$-CNG with an adaptive local pruning rule that avoids manually tuning the parameter $\alpha$ and preserves more useful shortcut edges. We also propose optimizations to accelerate graph construction further. Empirical results show $\alpha$-CNG consistently outperforms state-of-the-art PG methods, achieving superior accuracy-efficiency trade-offs. Furthermore, our edge pruning rule demonstrates transferability, enhancing query performance when integrated into other popular PGs.

### v5

@@ 7. Conclusions
    This paper introduces $\alpha$-CG, a new PG structure for high-performance ANN search. Specifically, $\alpha$-CG employs a well-designed pruning rule to eliminate ineffective candidates. We prove that, under a realistic assumption that the distance between the query and its NN is bounded by a constant $\tau$, $\alpha$-CG guarantees exact NN retrieval in poly-logarithmic time. Without this assumption, it ensures ANN search within the same complexity bounds. To reduce graph construction overhead, we develop an approximate variant $\alpha$-CNG with an adaptive local pruning rule that avoids manually tuning the parameter $\alpha$ and preserves more useful shortcut edges. We also propose optimizations to accelerate graph construction further. Empirical results show $\alpha$-CNG consistently outperforms state-of-the-art PG methods, achieving superior accuracy-efficiency trade-offs. Furthermore, our edge pruning rule demonstrates transferability, enhancing query performance when integrated into other popular PGs.
    Acknowledgements.
    Shangqi Lu’s research was supported by a start-up fund from HKUST-Guangzhou.